OpenStudy (anonymous):

Determine whether the graph of y = 3x2 − 7x + 2 opens up or down and whether it has a maximum or minimum point. think the answer is maximum am I correct ?

OpenStudy (wolf1728):

The value of the coefficient of x² determines the way the parabola will be aligned. If the coefficient of x² is positive: the parabola is concave up (shaped like a 'U') and because of the 'U' shape, the function has a minimum point.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[y=3\left( x ^{2}-\frac{ 7 }{ 3 } x+\left( \frac{ 7 }{ 6 } \right)^{2}-\left( \frac{ 7 }{ 6 } \right)^{2}\right)+2\] \[y=3\left( x-\frac{ 7 }{ 6 } \right)^{2}-3*\frac{ 49 }{36 }+2\] \[y+\frac{ 25 }{ 12 }=3\left( x-\frac{ 7 }{ 6 } \right)^{2}\] \[\left( x-\frac{ 7 }{6 } \right)^{2}=\frac{ 1 }{ 3 }\left( y+\frac{ 25 }{12 } \right)\] It is of the form \[X ^{2}=4 a Y\] It is an upward parabola with vertex \[\left( \frac{ 7 }{6 },\frac{ -25 }{12 } \right)\] this is the point of minima.